OpenStudy (anonymous):

which of the following is equivalent to log32^4 when i put it into my calculator it gives me 6.02 but my answers to choose from are .04 .0424 .125 2.5

OpenStudy (wolf1728):

32^4 = 1,048,576.00 and the log (base 10) of that is 6.02 so at least your calculator is correct.

OpenStudy (wolf1728):

I wonder if log 32^4 is meant for another number base? Perhaps base 4?

OpenStudy (wolf1728):

The base 4 log of 32 is 2.5

OpenStudy (anonymous):

ok so wait when i put that into a calculator how should i put it in?

OpenStudy (owlcoffee):

I was thinking the same, perhaps a minconception of the problem.

OpenStudy (wolf1728):

I'd input it as log (1,048,576)

OpenStudy (wolf1728):

Tiffani, How is the problem worded?

OpenStudy (owlcoffee):

use the definition of logarithm.

OpenStudy (wolf1728):

32 = 2^5 and 4 = 2^2 I wonder if that has anything to do with the solution.

OpenStudy (anonymous):

HI :D

OpenStudy (wolf1728):

Hello sourwing

OpenStudy (amoodarya):